# Linear Bandits with Partially Observable Features

### v1

@@ 4.2 Doubly Robust Lasso Estimator
8:         Compute $\widehat{a}_{t}:=\arg\max_{a\in[K]}{\widetilde{\mathbf{x}}_{a}}^{\top}\widehat{\boldsymbol{\mu}}^{L}_{t-1}$
9:         while $\widetilde{a}_{t}\neq a_{t}$ and $\text{count}\leq\rho_{t}$ do
10:              Sample $a_{t}$ with $\mathbb{P}(a_{t}=\widehat{a}_{t})=1-(t^{-1/2})$ and $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1),\;\forall k\neq\widehat{a}_{t}$.
11:              Sample $\widetilde{a}_{t}$ according to  Eq. 9.
12:              $\text{count}=\text{count}+1$
13:     Play $a_{t}$ and observe $y_{a_{t},t}$.
14:     if $\widetilde{a}_{t}\neq a_{t}$ then
15:         Set $\widehat{\boldsymbol{\mu}}^{L}_{t}:=\widehat{\boldsymbol{\mu}}^{L}_{t-1}$
16:     else
17:         Update $\widehat{\boldsymbol{\mu}}^{L}_{t}$ following Eq. 11 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}^{L}_{t}$ following Eq. 8.
5 Proposed Algorithm and Regret Analysis
In this section, we propose Algorithm 1, a novel estimation-based algorithm for handling partially observable features.
5.1 Robust to Latent Features (RoLF) Algorithm
In the initialization step, when the observed features are given, our algorithm finds a set of orthogonal basis $\{\mathbf{b}_{1},\dots,\mathbf{b}_{K-d}\}\in\mathrm{R}(\mathbf{X})^{\perp}$ to augment each observed features. After the forced-exploration phase, the algorithm computes the candidate action, denoted by $\widehat{a}_{t}$, and then resample both $\widetilde{a}_{t}$ and $a_{t}$ until they match. When the resampling phase ends, and the agent selects $a_{t}$ and observes $y_{a_{t},t}$. If they match within $\rho_{t}:=\log((t+1)^{2}/\delta^{\prime})/\log(1/p)$, the algorithm updates both the imputation and main estimators; otherwise neither estimator is updated.
The proposed algorithm does not require the knowledge of the dimension of the latent features $d_{u}$ and the dimension of the projected rewards from latent feature space onto the $\mathrm{R}(\mathbf{X})^{\perp}$. Although we present the algorithm on fixed feature vectors, the algorithm applies to arbitrary feature vectors that changes over time by updating the orthogonal basis.
5.2 Regret Analysis
We analyze the regret bound of RoLF using Lasso estimators, as stated in the following theorem:
Theorem 3 (Regret bound for Lasso RoLF).
Let $d_{h}$ denote the dimension of the projected latent rewards defined in Eq. 6. Then for $\delta\in(0,1)$ and $p\in(1/2,1)$, the expected cumulative regret of the proposed algorithm is bounded by
$\begin{split}\mathbb{E}[\mathrm{Reg}(T)]&\leq 10\delta\log T+\frac{16K^{2}(\sqrt{K}+p^{-1})^{2}}{(1-p)^{2}}+\frac{4p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+\frac{80\sigma}{p}\sqrt{2(d+d_{h})T\log\frac{2KT^{2}}{\delta}},\end{split}$ (13)
To the best of our knowledge, Theorem 3 is the first regret bound sublinear in $T$ for the latent features without any structural assumption. Note that the number of rounds for the exploration phase is $O(K^{3}\log KT)$, which is only logarithmic in the horizon $T$. The factor $K^{3}$ is not reducible since the algorithm must estimate all $K$ biases from the missing features. Using the Gram matrix with full feature vectors, $\sum_{a=1}^{K}\widetilde{\mathbf{x}}_{a}\widetilde{\mathbf{x}}_{a}^{\top}$ in combination with DR estimation reduces the exploration phase time from $O(K^{4}\log KT)$ to $O(K^{3}\log KT)$, reducing the complexity by a factor of $K$. The convergence rate in the last term is proportional to $\sqrt{d+d_{h}}$ rather than $\sqrt{K}$, as shown in Eq. 12. Thus, our regret bound is $O(\sqrt{(d+d_{h})T\log KT})$.
6 Numerical Experiments
In this experiment, we simulate and compare two versions of our algorithm, presented in Algorithm 1 and Algorithm 2 (Appendix A), with linear bandit algorithms that use only observed features: LinUCB (Li et al., 2010; Chu et al., 2011) and LinTS. These algorithms use the UCB and Thompson sampling methods, respectively, when the reward is modeled as a linear function of the features. Additionally, since our algorithm incorporates DR estimation with the Lasso estimator, we include DRLasso (Kim and Paik, 2019) in the comparison as well. To further evaluate the performance of our algorithm in scenarios where latent features are expected but ignored, we also compare it with UCB($\delta$) (Lattimore and Szepesvári, 2020), an MAB algorithm without features.
For the simulation environment, we generate true features $\mathbf{z}_{a}$ for each arm $a\in[K]$ from $\mathcal{N}(\mathbf{0},\mathbf{I}_{d_{z}})$ and subsample $d$ elements to construct $\mathbf{x}_{a}$. Orthogonal basis vectors $\{\mathbf{b}_{1},\dots,\mathbf{b}_{K-d}\}$ are derived via singular value decomposition (SVD) on the observed feature matrix $\mathbf{X}$, ensuring orthogonality to $\mathrm{R}(\mathbf{X})$. We augment $\mathbf{X}$ with the basis vectors via linear concatenation. Rewards are generated by sampling the unknown parameter $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{k}$ from $\mathrm{Unif}(-1/2,1/2)$. The hyperparameter $p$, for the sampling distribution of $\widetilde{a}_{t}$, is set to $0.6$ (see Eq. 9). The confidence parameter $\delta$ is $10^{-4}$, and the total decision horizon is $T=1200$. To address both partial and full observability, $d_{z}\geq d$ is used, and we run 5 independent experiments. We compare the algorithms across three scenarios:
(a) $d=1$
(b) $d=\left\lfloor\frac{d_{z}}{2}\right\rfloor$
(c) $d=d_{z}-1$
Figure 2: Cumulative regrets of the algorithms in comparison for scenario 1 ($K=50,d_{z}=31$).
(a) $K=d=20$
(b) $K=d=30$
(c) $K=d=40$
Figure 3: Cumulative regrets of the algorithms in comparison for scenario 2 ($d_{z}=60$).
(a) $K=15$, $d=30$
(b) $K=20$, $d=40$
(c) $K=30$, $d=60$
Figure 4: Cumulative regrets of the algorithms in comparison for scenario 3 ($d_{z}=d$).
Scenario (i).
We examine algorithm performance as $d$, the number of observed elements, varies to assess the impact of observability. With $K=50$ arms and $d_{z}=31$, we compare results for $d=1$, $\lfloor d_{z}/2\rfloor=15$, and $d_{z}-1=30$. Figure 2 presents the results, showing that our algorithm consistently outperforms others in regret and robustness. In contrast, LinUCB, LinTS, and DRLasso show significant dependence on the number of observed features, with performance deteriorating and variability increasing as observability decreases.
Scenario (ii).
In this scenario, the number of arms is equal to the dimension of the observed features, i.e., $K=d$. The main objective of this experiment is to demonstrate that our algorithm remains robust to changes in the number of arms $K$, unlike MAB algorithms that disregard the observed features. Specifically, we compare the algorithms with $K$ set to 20, 30, and 40, and $d_{z}=60$ for all cases. Figure 3 shows the results for this scenario. As we can observe, the performance of UCB($\delta$) worsens as the number of arms increases in each environment. In contrast, our algorithm shows better performance in terms of both the level of regret and robustness.
Scenario (iii).
We evaluate performance when the number of arms is less than the dimension of observed features, setting $d=2K$ and varying $K$ as 15, 20, and 30, with $d_{z}=d$. Before using the features in our algorithms, we apply singular value decomposition (SVD) for dimensionality reduction. Figure 4 shows that our algorithm performs well even in extreme cases. By applying dimension reduction through SVD, our algorithm remains applicable even when the matrix of feature vectors is not full rank. Furthermore, the results suggest that our algorithm demonstrates superior performance even in the absence of partial observability.
7 Conclusion
In this paper, we addressed the challenges due to partially observable features in linear bandit problems. We showed that algorithms relying solely on observed features may suffer linear regret due to model misspecification and introduced RoLF, a novel algorithm that incorporates latent features without prior knowledge of them. Our algorithm improves the regret bound over traditional methods that ignore latent features, with numerical experiments supporting the theoretical results. Moving forward, several open directions remain for future work. First, while the linear bandit framework allows our feature augmentation strategy in Section 4.1 to be interpreted as another linear bandit problem, extending this approach to more general reward functions, such as generalized linear models, would be interesting. Furthermore, as noted in Section 3.2, we assumed the latent portion of reward to be linear with respect to the unobserved features. However, the unobserved reward can be viewed as an exogenous factor affecting the agent’s learning. This perspective enables extending the unobserved reward to a general function class without structural assumptions.
Appendix A Robust to Latent Feature Algorithm with Ridge Estimator
1:INPUT: features $\{\mathbf{x}_{a}:a\in[K]\}$, coupling probability $p\in(1/2,1)$, confidence parameter $\delta>0$.
2:Initialize $\widehat{\boldsymbol{\mu}}_{0}=\mathbf{0}_{K}$, the exploration phase $\mathcal{E}_{t}=\emptyset$ and the exploration factor $C_{\mathrm{e}}:=32(1-p)^{-2}K^{2}$.
3:Find orthogonal basis $\mathbf{b}_{1},\ldots,\mathbf{b}_{K-d}$ in $\mathrm{R}(\mathbf{X})^{\perp}$ to construct $\{{\widetilde{\mathbf{x}}_{a}}:a\in[K]\}$
4:for $t=1,\ldots,T$ do
5:     if $|\mathcal{E}_{t}|\leq C_{\mathrm{e}}\log(2Kt^{2}/\delta)$ then
7:     else
8:         Compute $\widehat{a}_{t}:=\arg\max_{a\in[K]}{\widetilde{\mathbf{x}}_{a}}^{\top}\widehat{\boldsymbol{\mu}}^{R}_{t-1}$
9:         while $\widetilde{a}_{t}\neq a_{t}$ and $\text{count}\leq\rho_{t}$ do
10:              Sample $a_{t}$ with $\mathbb{P}(a_{t}=\widehat{a}_{t})=1-(t^{-1/2})$ and $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1),\;\forall k\neq\widehat{a}_{t}$.
11:              Sample $\widetilde{a}_{t}$ according to  Eq. 9.
12:              $\text{count}=\text{count}+1$
13:     Play $a_{t}$ and observe $y_{a_{t},t}$.
14:     if $\widetilde{a}_{t}\neq a_{t}$ then
15:         Set $\widehat{\boldsymbol{\mu}}^{R}_{t}:=\widehat{\boldsymbol{\mu}}^{R}_{t-1}$
16:     else
17:         Update $\widehat{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 14 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 15.
@@ Appendix A Robust to Latent Feature Algorithm with Ridge Estimator
For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. Then for all round $t$ such that $|\mathcal{E}_{t}|\geq 32(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-3\delta$,
@@
For $\delta\in(0,1)$, the expected cumulative regret of the proposed algorithm using DR Ridge estimator is bounded by
$\mathrm{Reg}(T)\leq 6\delta\log T+\frac{2p\sqrt{T}}{K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{T^{2}}{\delta}}+1\right).$
@@
Appendix B A Modified Algorithm for Time-Varying Observed Features
In this section, we propose an algorithm for linear bandits with partially observable features under the setting where the observed features vary over time.
B.1 Problem Formulation
@@
B.2 Proposed Method: Orthogonal Basis Augmentation
We address the problem by augmenting Euclidean basis $\mathbf{e}_{1},\ldots,\mathbf{e}_{K}$ in $\mathbb{R}^{K}$ to estimate bias caused by the unobserved features. Let $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{K}]\in\mathbb{R}^{d+K}$ and let $\Delta_{a}:=\langle\mathbf{u}_{a},\theta_{\star}^{(u)}\rangle$ denote the bias stems from the latent features. Then,
$y_{a,t}$ $=\langle\mathbf{x}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\rangle+\epsilon_{a,t}$
$=\langle\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\mathbf{e}_{1}\cdots\mathbf{e}_{K}],[\boldsymbol{\theta}_{\star}^{(o)}\Delta_{1}\cdots\Delta_{K}]\rangle+\epsilon_{a,t}.$
@@ B.2 Proposed Method: Orthogonal Basis Augmentation
If observed features are vary over time, for $\delta\in(0,1)$, the expected cumulative regret of the proposed algorithm Ridge-RoLF-V using DR Ridge estimator is bounded by
$\begin{split}\mathrm{Reg}(T)\leq&6\delta\log T+\frac{2p\sqrt{T}}{d+K-1}\frac{\log\frac{(T+1)^{2}}{\delta}}{\log(1/p)}+\frac{32(K+d)^{2}}{(1-p)^{2}}\log\frac{2(K+d)T^{2}}{\delta}\\ &+8\sqrt{(d+K)T}\left(\frac{\sigma}{p}\sqrt{\log\frac{T^{2}}{\delta}}+1\right).\end{split}$
@@
Appendix C Missing Proofs
C.1 Proof of Theorem 1
@@
For brevity, we denote the latent reward components as $g_{1}:=u_{1}\theta_{\star}^{(u)}$ and $g_{2}:=u_{2}\theta_{\star}^{(u)}$, yielding that $g_{1}=-3$ and $g_{2}=-19/4$. Note that since $\gamma_{2}\neq 2\gamma_{1}$ and $|{g_{i}}|\geq 3>0$ for all $i\in\{1,2\}$, thus our problem setup satisfies the “large deviation” criterion in Definition 1 and Theorem 2 of Ghosh et al. [2017], by letting $l=3$ and $\beta=0$. Consequently, by applying the theorem, it follows that using OFUL in this problem instance results in linear regret with respect to $T$, i.e., $\Omega(T)$. Inspired by this theorem, we demonstrate that LinTS encounters the same issue.
@@
$\begin{split}\hat{\boldsymbol{\theta}}_{t}&=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}\mathbf{Y}_{t})\\ &=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}(\mathbf{X}_{t}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{g}_{t}+\boldsymbol{\epsilon}_{t}))\\ &=\boldsymbol{\theta}_{\star}^{(o)}-\lambda\mathbf{V}_{t}^{-1}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\mathbf{g}_{t}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\boldsymbol{\epsilon}_{t},\end{split}$ (16)
@@
$\hat{\theta}_{t}=\theta_{\star}^{(o)}-\frac{\theta_{\star}^{(o)}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1},$
where we assume $\lambda=1$. Note also that we denote $\hat{\boldsymbol{\theta}}_{t}$ and $\boldsymbol{\theta}_{\star}^{(o)}$ by $\hat{\theta}_{t}$ and $\theta_{\star}^{(o)}$, respectively, since both are scalars. Hence, the estimation error is computed as:
$\hat{\theta}_{t}-\theta_{\star}^{(o)}=-\frac{\theta_{\star}^{(o)}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}.$ (17)
@@
$\geq\bar{g}N_{1}+2\bar{g}N_{2}$
$=\bar{g}N_{1}+2\bar{g}(t-N_{1})$
$=2\bar{g}t-\bar{g}N_{1}$
$\geq\bar{g}t\qquad\qquad(\because N_{1}\leq t)$
@@
where $\bar{g}=\min\{g_{1},g_{2}\}=-19/4$, which implies that $\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}=\Theta(t)$. For the denominator, $\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1$, as it grows at a rate of $O(t)$, implying that the second term in Eq. 17 is $\Theta(1)$, and that $\hat{\theta}_{t}$ is not consistent since it does not converge to $\theta_{\star}^{(o)}$ as $t\to\infty$.
For arm 2, which is optimal, to be selected in round $t+1$ under LinTS, the condition $x_{2}\widetilde{\theta}_{t}\geq x_{1}\widetilde{\theta}_{t}$ must hold, where $\widetilde{\theta}_{t}\sim\mathcal{N}\left(\hat{\theta}_{t},\frac{v^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}\right)$. Given the assumptions that $x_{1}=1$ and $x_{2}=2$, arm 2 is selected whenever $\widetilde{\theta}_{t}\geq 0$. Thus, for arm 1 to be chosen, we require $\widetilde{\theta}_{t}<0$. We will show that the probability of $\widetilde{\theta}_{t}<0$ does not diminish sufficiently to be ignored even the agent plays sufficiently large amount of time. To clarify, let us define two events $E_{\widetilde{\theta}}:=\{\widetilde{\theta}_{t}\geq 0\}$ and $E_{\hat{\theta}}:=\{\hat{\theta}_{t}\geq 0\}$.
We revisit Eq. 16 as follows:
$\hat{\theta}_{t}$ $=\frac{\theta_{\star}^{(o)}\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}$
@@
where for the second term of Eq. 18, for $t\geq 20$, it is upper bounded by:
@@
$\leq-\frac{19}{8}+\frac{19}{8t}$
$\leq-\frac{9}{4},$
resulting in
$\begin{split}\hat{\theta}_{t}&\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq 2-\frac{9}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &=-\frac{1}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}\end{split}$ (19)
@@
$\mathbb{P}(\hat{\theta}_{t}>0)\leq\mathbb{P}\left(-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>0\right)=\mathbb{P}\left(\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{4}\right).$
@@
$\mathbb{P}(\hat{\theta}_{t}>0)\leq\exp\left(-\frac{t}{128\sigma^{2}}\right).$

### v2

@@ Appendix C Robust to Latent Feature Algorithm with Ridge Estimator
8:        Compute $\widehat{a}_{t}:=\arg\max_{a\in[K]}\widetilde{\mathbf{x}}_{a}^{\top}\widehat{\boldsymbol{\mu}}^{R}_{t-1}$.
9:     end if
10:     while $\widetilde{a}_{t}\neq a_{t}$ and $\text{count}\leq\rho_{t}$ do
11:        Sample $a_{t}$ with $\mathbb{P}(a_{t}=\widehat{a}_{t})=1-(t^{-1/2})$ and $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1),\;\forall k\neq\widehat{a}_{t}$.
12:        Sample $\widetilde{a}_{t}$ according to  Eq. 9.
13:        $\text{count}=\text{count}+1$.
14:     end while
15:     Play $a_{t}$ and observe $y_{a_{t},t}$.
16:     if $\widetilde{a}_{t}\neq a_{t}$ then
17:        Set $\widehat{\boldsymbol{\mu}}^{R}_{t}:=\widehat{\boldsymbol{\mu}}^{R}_{t-1}$.
18:     else
19:        Update $\widehat{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 12 with $\widetilde{y}_{a,t}$ and update $\check{\boldsymbol{\mu}}_{t}^{R}$ following Eq. 13.
20:     end if
21:  end for
@@
For each $t$, let $\mathcal{E}_{t}\subseteq[t]$ denote an exploration phase such that for $\tau\in\mathcal{E}_{t}$ the action $a_{\tau}$ is sampled uniformly over $[K]$. Assume that $\|\boldsymbol{\mu}_{\star}\|_{\infty}\leq 1$ and $\|\widetilde{\mathbf{x}}_{a}\|_{\infty}\leq 1$ for all $a\in[K]$. Then for all round $t$ such that $|\mathcal{E}_{t}|\geq 32(1-p)^{-2}K^{2}\log(2Kt^{2}/\delta)$, with probability at least $1-3\delta$,
@@
Suppose $\|\boldsymbol{\mu}_{\star}\|_{\infty}\leq 1$ and $\|\mathbf{z}_{a}\|_{\infty}\leq 1$ for all $a\in[K]$. For $\delta\in(0,1)$, with probability at least $1-4\delta$, the cumulative regret of the proposed algorithm using DR Ridge estimator is bounded by
$\mathrm{Reg}(T)\leq\frac{32K^{2}}{(1-p)^{2}}\log\frac{2dT^{2}}{\delta}+4\sqrt{2T\log\frac{2}{\delta}}+\frac{4\sqrt{T}}{K-1}+4\delta+8\sqrt{KT}\left(\frac{\sigma}{p}\sqrt{\log\frac{t+1}{\delta}}+1\right).$
@@
Appendix D A Modified Algorithm for Time-Varying Observed Features
In this section, we define the problem of linear bandits with partially observable features under the setting where the observed features vary over time, describe our proposed method, and provide theoretical guarantees.
D.1 Problem Formulation
@@
D.2 Proposed Method: Orthogonal Basis Augmentation
We address the problem by augmenting standard basis $\mathbf{e}_{1},\ldots,\mathbf{e}_{K}$ in $\mathbb{R}^{K}$ to estimate bias caused by the unobserved features. Let $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{K}]\in\mathbb{R}^{d+K}$ and let $\Delta_{a}:=\langle\mathbf{u}_{a},\theta_{\star}^{(u)}\rangle$ denote the bias stems from the latent features. Then,
$y_{a,t}$ $=\langle\mathbf{x}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\rangle+\epsilon_{t}$
$=\langle\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\mathbf{e}_{1}\cdots\mathbf{e}_{K}],[\boldsymbol{\theta}_{\star}^{(o)}\Delta_{1}\cdots\Delta_{K}]\rangle+\epsilon_{t}.$
@@ D.2 Proposed Method: Orthogonal Basis Augmentation
If observed features are vary over time, for $\delta\in(0,1)$, with probability at least $1-4\delta$, the cumulative regret of the proposed algorithm Ridge-RoLF-V using DR Ridge estimator is bounded by
$\begin{split}\mathrm{Reg}(T)\leq&4\delta+\frac{2\sqrt{T}}{d+K-1}+\frac{32(K+d)^{2}}{(1-p)^{2}}\log\frac{2(K+d)T^{2}}{\delta}+8\sqrt{(d+K)T}\left(\frac{\sigma}{p}\sqrt{\log\frac{T^{2}}{\delta}}+1\right).\end{split}$
@@
Appendix E Missing Proofs
E.1 Proof of Theorem 1
@@
For brevity, we denote the latent reward components as $g_{1}:=u_{1}\theta_{\star}^{(u)}$ and $g_{2}:=u_{2}\theta_{\star}^{(u)}$, yielding that $g_{1}=-3$ and $g_{2}=-19/4$. Note that since $\gamma_{2}\neq 2\gamma_{1}$ and $|{g_{i}}|\geq 3>0$ for all $i\in\{1,2\}$, thus our problem setup satisfies the “large deviation” criterion in Definition 1 and Theorem 2 of Ghosh et al. (2017), by letting $l=3$ and $\beta=0$. Consequently, by applying the theorem, it follows that using OFUL in this problem instance results in linear regret with respect to $T$, i.e., $\Omega(T)$. Inspired by this theorem, we demonstrate that LinTS encounters the same issue.
@@
$\begin{split}\widehat{\boldsymbol{\theta}}_{t}&=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}\mathbf{Y}_{t})\\ &=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}(\mathbf{X}_{t}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{g}_{t}+\boldsymbol{\epsilon}_{t}))\\ &=\boldsymbol{\theta}_{\star}^{(o)}-\lambda\mathbf{V}_{t}^{-1}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\mathbf{g}_{t}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\boldsymbol{\epsilon}_{t},\end{split}$ (14)
@@
$\widehat{\theta}_{t}=\theta_{\star}^{(o)}-\frac{\theta_{\star}^{(o)}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1},$
where we assume $\lambda=1$. Note also that we denote $\widehat{\boldsymbol{\theta}}_{t}$ and $\boldsymbol{\theta}_{\star}^{(o)}$ by $\widehat{\theta}_{t}$ and $\theta_{\star}^{(o)}$, respectively, since both are scalars. Hence, the estimation error is computed as:
$\widehat{\theta}_{t}-\theta_{\star}^{(o)}=-\frac{\theta_{\star}^{(o)}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}.$ (15)
@@
$\geq\underline{g}N_{1}+2\underline{g}N_{2}$
$=\underline{g}N_{1}+2\underline{g}(t-N_{1})$
$=2\underline{g}t-\underline{g}N_{1}$
$\geq\underline{g}t\qquad\qquad(\because N_{1}\leq t)$
@@
where $\underline{g}=\min\{g_{1},g_{2}\}=-19/4$, which implies that $\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}=\Theta(t)$. For the denominator, $\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1$, as it grows at a rate of $O(t)$, implying that the second term of the right-hand side in Eq. 15 is $\Theta(1)$, and that $\widehat{\theta}_{t}$ is not consistent since it does not converge to $\theta_{\star}^{(o)}$ as $t\to\infty$.
For arm 2, which is optimal, to be selected in round $t+1$ under LinTS, the condition $x_{2}\widetilde{\theta}_{t}\geq x_{1}\widetilde{\theta}_{t}$ must hold, where $\widetilde{\theta}_{t}\sim\mathcal{N}\left(\widehat{\theta}_{t},\frac{v^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}\right)$. Given the assumptions that $x_{1}=1$ and $x_{2}=2$, arm 2 is selected whenever $\widetilde{\theta}_{t}\geq 0$. Thus, for arm 1 to be chosen, we require $\widetilde{\theta}_{t}<0$. We will show that the probability of $\widetilde{\theta}_{t}<0$ does not diminish sufficiently to be ignored even the agent plays sufficiently large amount of time. To clarify, let us define two events $E_{\widetilde{\theta}}:=\{\widetilde{\theta}_{t}\geq 0\}$ and $E_{\hat{\theta}}:=\{\widehat{\theta}_{t}\geq 0\}$. We revisit Eq. 14 as follows:
$\widehat{\theta}_{t}$ $=\frac{\theta_{\star}^{(o)}\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}+\frac{\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}$
@@
where for the second term of Eq. 16, for $t\geq 19$,
@@
$\leq-\frac{19}{8}+\frac{19}{8t},$
which is upper bounded by $-{9}/{4}$. This bound is followed by:
$\begin{split}\widehat{\theta}_{t}&\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq 2-\frac{9}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &=-\frac{1}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}.\end{split}$ (17)
@@
$\mathbb{P}(\widehat{\theta}_{t}>0)\leq\mathbb{P}\left(-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>0\right)=\mathbb{P}\left(\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}>\frac{1}{4}\right).$
@@
$\mathbb{P}(\widehat{\theta}_{t}>0)\leq\exp\left(-\frac{t}{128\sigma^{2}}\right).$